Mathematicians have been searching for a formula that yields prime numbers. One such formula was x2-x+41. Select some numbers for x, substitute them in the formula, and see if prime numbers occur. Try to find a number for x that when substituted in the formula yields a composite number. What number will solve for x?
Answer by themathtutor2009(81)

x=1: 1^2-1+41=41....prime

x=5: 5^2-5+41=61....prime

x=10: 10^2-10+41=131... prime

x=20: 20^2-20+41=421...prime

x=40: 40^2-40+41=1601...prime

x=41: 41^2-41+41=1681...composite

x=42: 42^2-42+41=1763...composite

x=45: 45^2-45+41=2021...composite

Any non-negative number less than 41 yields a prime number